OpenStudy (anonymous):

A sportswear store sells baseball caps with the local baseball team's logo on them. Last year, the store sold 600 caps at $15 each. The store manager is planning to increase the price. A consumer survey shows that for every $1 increase, they will be a drop of 30 sales a year.

OpenStudy (anonymous):

a) What should the selling price be to maximize the annual revenue? b) What is the maximum annual revenue from the caps?

OpenStudy (anonymous):

The answer for a) is supposed to be: 84.5 and b) 6.5, 6.5. How?

OpenStudy (anonymous):

if the answer for A is 84.5, it is against on the survey. .

OpenStudy (anonymous):

given that the price per cap is 15, and survey states that their will be a drop of sales by 30 number of caps if they will increase 1 per cap, how could the answer be 84.5? i think the answer for the increase is less than 1 dollar. that will be 15+x; where x is less than 1 that's my interpretation to the problem. .

sam (.sam.):

Let R be revenue per year, in dollars. R = units sold multiply price per unit. Let units sold be U caps. Let price per unit be P, in dollars per cap. Therefore: R = UP. to begin with, P = $15 per cap, and U = 600 caps. To increase the price by $1, and drop sales by 30: R = (600 - 30) (15 + 1). To do the same thing again: R = (600 - 30 - 30) (15 + 1 + 1) = (600 - 60)(15 + 2). Notice that these will decrease by multiples of 30 and increase by multiples of 1 respectively. Give the number of repetitions the value n. So: U = 600 - 30n. P = 15 + n. So: R = UP = (600 - 30n) (15 + n). Expanding: R = 9000 + 600n - 450n - 30n^2. R = 9000 + 150n - 30n^2. Factorising: R = -30 [n^2 - 5n - 30 ]. Completing the square: R = -30 [ (n - 2½)^2 - (2½^2) - 30 ]. R = -30 [ (n - 2½)^2 - 36¼ ]. R = (30)36¼ - 30(n - 2½)^2. Since you have completed the square, any derivative will only be 0 at the point when the square is 0. i.e. (n - 2½)^2 must equal 0. (n - 2½)^2 = 0. Square rooting: n - 2½ = 0. n = 2½. Differentiating: dR/dn = -60(n - 2½) = 150 - 60n. Solving for 0: dR/dn = 0 → 150 - 60n = 0. 150 = 60n. n = 15/6 = 2½. So, for (a) : Let n = 2½. P = 15 + n = 15 + 2½ = 17½. Thus P = $17.50 per cap. For (b) : Again, n = 2½. R = (600 - 30n) (17½). R = (600 - 30(2½)) (17½). R = (600 - 75) (17½). R = (525)17½. R = $9187.50 per year.

OpenStudy (anonymous):

woah. That's how to get the answer?